# A multi-task deep learning approach for lane-level pavement performance prediction with segment-level data

### v1

@@ Abstract
The elaborate pavement performance prediction is an important premise of implementing preventive maintenance. Our survey reveals that in practice, the pavement performance is usually measured at segment-level, where an unique performance value is obtained for all lanes within one segment of 1km length. It still lacks more elaborate performance analysis at lane-level due to costly data collection and difficulty in prediction modeling. Therefore, this study developed a multi-task deep learning approach to predict the lane-level pavement performance with a large amount of historical segment-level performance measurement data. The unified prediction framework can effectively address inherent correlation and differences across lanes. In specific, the prediction framework firstly employed an Long Short-Term Memory (LSTM) layer to capture the segment-level pavement deterioration pattern. Then multiple task-specific LSTM layers were designed based on number of lanes to capture lane-level differences in pavement performance. Finally, we concatenated multiple task-specific LSTM outputs with auxiliary features for further training and obtained the lane-level predictions after fully connected layer. The aforementioned prediction framework was validated with a real case in China. It revealed a better model performance regardless of one-way 2-lane, 3-lane, and 4-lane scenarios, all lower than 10% in terms of mean absolute percentage error. The proposed prediction framework also outperforms other ensemble learning and shallow machine learning methods in almost every lane.
@@
journal: Automation in Construction
@@
[inst1]organization=School of Transportation,addressline=Southeast University, city=Nanjing, postcode=210000, state=Jiangsu, country=China
@@
A multi-task deep learning approach is proposed for lane-level pavement performance prediction.
The unified prediction framework can effectively address inherent correlation and differences across lanes.
Pavement performance indicator time series is processed using shared LSTM and task-specific head and get a initial predictions.
Auxiliary pavement features is concatenated with initial predictions to adjust the predictions and get the final predictions.
The model is validated on one-way 2-lane, 3-lane, and 4-lane scenarios, all lower than 10% in terms of mean absolute percentage error.
@@ 1 Introduction
The most widely used pavement performance measurements include pavement condition index (PCI), pavement quality index (PQI), and riding quality index (RQI). These measurements in a lane with shorter length can assist in identifying small but severely damaged road sections that are detrimental to traffic. The performance measurement prediction can further provide refined performance trends in future years for each lane. Utilizing the predicted lane-level performance values can further execute the preventive maintenance in a precisely and cost-effective way. Thus, lane-level performance predictions are significantly more useful for pavement maintenance than segment-level ones that merely provide one rough value for all lanes in a segment of 1 km. However, it is impossible of collecting huge amount of lane-level data in reality and directly predicting the performance in next few years. To address such difficulty, this study is dedicated to developing a lane-level pavement performance prediction framework with a large amount of segment-level data.
@@
To address the difficulties during lane-level performance prediction, we propose a multi-task deep learning approach for lane-level pavement performance prediction based on segment-level data to deal with those challenges. The model employs an LSTM-based shared layer to capture the common and representative changing pattern information from pavement performance time series. Based on the shared information, the task-specific heads are used to extract the details and changes related to a specific lane and generate predictions for the lane. Subsequently, the predictions are horizontally concatenated with auxiliary features, and the features distinguish roads with different decay patterns. At last, the task-specific output layer will handle the concatenation results to get the final lane-level predictions. To verify the proposed model, the real case was employed to check prediction accuracy of PCI, PQI, and RQI on three scenarios including one-way 2-lane, 3-lane, and 4-lane roads.
The remainder of this paper is organized as follows: In section 2, we summarize the related work on pavement performance prediction. In section 3, we present the structure of the proposed model. Then, we validate the model through comparative and ablation experiments based on a dataset collected on our own in section 4. Finally, we present our conclusions and offer guidelines for future work.
@@ 2 RELATED WORK
Pavement performance prediction models can be categorized into two types: Non-data-driven and data-driving models. The former rely on specific mathematical equations or physical principles. Still, the latter primarily relies on a large amount of data input rather than solely on theoretical assumptions or prior knowledge.
2.1 Non-data-driven models
@@
Non-data-driven models predict pavement performance based on deterministic principles and equations, like empirical formula models and mechanical models. They can also take into account random factors in pavement performance deterioration and use probability and statistics methods, such as Markov models and Monte Carlo simulations, to describe the possible distribution or changing trend of pavement performance. Wu compared the effectiveness of pavement performance prediction based on S-shaped model, polynomial model, and exponential model. And the S-shaped model demonstrates the best predictive performance[3]. Cao selected service life and cumulative number of axle loads as independent variables and applied cosine deterioration equation to predict PCI and rutting depth index of roads with excellent pavement performance[4]. Sidess combining empirical-mechanistic approach and regressive empirical approach to predicting the international roughness index[5].Wang calibrated parameters of the rutting prediction model for semi-rigid base asphalt pavement in MEPDG based on real data From Xi’an, China[6].Dong predicted asphalt pavement performance utilizing Markov models based on insufficient inspection data[7]. Wang put forward a Grey–Markov combination model to predict PCI[8]. Since pavement deterioration rate may not be constant, Abaza developed a non-homogeneous Markov chain, which uses different transition matrix to predict future pavement performance for every time step[9]. Besides, Wasiq proposed a hybrid approach combining both homogeneous and non-homogeneous Markov Chain to predict the PCI[10].
Non-data-driven models are not suitable for considering the complex relationships between various factors and the coupled effects of some factors on pavement performance degradation.
2.2 Data-driven models
data-driven Models do not rely on specific mathematical equations or physical principles, but rather learn patterns in the data through algorithms and are used for the prediction of pavement performance. Wang developed a hybrid grey relation analysis and support vector machine regression model to predict pavement performance[11]. Xiao combined particle swarm optimization algorithm and back propagation neural network to predict asphalt pavement performance[12]. Tao proposed a differential evolution particle swarm optimization back propagation neural network to predict PCI[13]. Abdualmtalab investigated the combined effect of pavement distress on flexible pavement performance in two climate regions (wet freeze and wet no freeze) separated using multiple linear regression and artificial neural network[14]. Cai proposed a Causal-Temporal Graph Convolution Network, combining the Graph Convolution Networks) and LSTMs to capture causal features and temporal features simultaneously, to predict pavement performance[15]. Bukharin proposed a two-stage model combining LSTM and artificial neural network (ANN). The LSTM is used to learn the pavement deterioration pattern based on sequential data and ANN further learns the factors influencing deterioration to adjust the final predictions[16]. Guo proposed a weighted multi-output neural network to predict IRI, faulting, longitudinal crack and transverse crack simultaneously[17].
Data-driven models are more suitable for performance prediction of a large number of road segments, and they can capture different decay patterns of the road segments as well as the relationship between segments’ decay and influencing factors. Also it is suitable for extracting the similarity and difference of lanes’ decay patterns.
3 METHODS
@@ 3.1 Problem definition
This paper is dedicated to proposing a lane-level pavement performance prediction model. The prediction is based on segment-level pavement inspection data, and static factors influencing pavement deterioration.
$\hat{y}_{t+1,N}=f(y_{t-k:t},x_{t-k:t},S)$ (1)
where:
$\hat{y}_{t+1,N}$ is the predicted performance value of each lane,
$N$ is the number of lanes,
$y_{t-k:t}$ is the predictive indicator’s time series from time step $t-k$ to time step $t$, like PCI, PQI and RQI,
$x_{t-k:t}$ are the covariates’ time series from time step $t-k$ to time step $t$,
$S$ is the static covariates in the observational process that are invariant. For instance, pavement material,
$f(\cdot)$ is the relationship between the predictive indicator and covariates.
@@ 3.2 Model overview
In this section, we focus on introducing the structure of the proposed multi-task deep learning model (MTL), as shown in Figure 2. The LSTM-based shared layer, module (a) in Figure 2, will extract general decay patterns of different lanes from the pavement performance time series. Then, the output of the layer will be sent to module (b); the module consists of an LSTM and a fully connected layer (FC layer in Figure 2). The LSTM will predict based on the output, and the fully connected layer will encode the predictions. In module (c), the encoded predictions will be horizontally concatenated with auxiliary features. The concatenation result will be sent to each fully connected layer followed. The fully connected layer will adjust the prediction using auxiliary features and then get the final lane-level prediction results.
@@ 3.2.1 LSTM-based shared layer
The LSTM-based shared layer is initially used to handle the pavement performance time series. It is responsible for discovering and extracting common decay patterns of different lanes. LSTM is a type of recurrent neural network architecture designed to handle the vanishing and exploding gradient problems often encountered in traditional recurrent neural networks when dealing with long data sequences. LSTM units have a more complex internal structure than standard recurrent neural network cells. They consist of three main gates: the input gate, the forget gate, and the output gate. The input gate decides how much new information to let into the cell state, the forget gate determines which parts of the previous cell state to discard, and the output gate controls how much of the cell state is output as the hidden state. This gated mechanism allows LSTMs to selectively remember or forget information over long intervals, making them well-suited for time series prediction tasks.
(1) Forget gate:
The forget gate controls what information is removed from the memory cell. Two inputs $x_{t}$ (input at the particular time $t$) and $h_{t-1}$ (previous cell output) are fed to the gate and multiplied with weight matrices followed by the addition of bias. The resultant is passed through an activation function which gives a binary output. If for a particular cell state, the output is 0, the piece of information is forgotten and for output 1, the information is retained for future use.
$f_{t}=\sigma(W_{f}[h_{t-1},x_{t}]+b_{f})$ (2)
(2) Input gate:
The input gate controls what information is added to the memory cell. First, the information is regulated using the sigmoid function and filter the values to be remembered similar to the forget gate using inputs $h_{t-1}$ and $x_{t}$. Then, a vector is created using tanh function that gives an output from -1 to +1, which contains all the possible values from $h_{t-1}$ and $x_{t}$. At last, the values of the vector and the regulated values are multiplied to obtain the useful information. The equation for the input gate is:
$i_{t}=\sigma(W_{i}[h_{t-1},x_{t}]+b_{i})$ (3)
@@
We multiply the previous state by $f_{t}$, disregarding the information we had previously chosen to ignore. Next, we include $i_{t}$*$C_{t}$. This represents the updated candidate values, adjusted for the amount that we chose to update each state value.
$c_{t}=f_{t}\odot c_{t-1}+i_{t}\odot\stackrel{{\scriptstyle\sim}}{{c_{t}}}$ (5)
@@
The output gate controls what information is output from the memory cell. First, a vector is generated by applying tanh function on the cell. Then, the information is regulated using the sigmoid function and filter by the values to be remembered using inputs $h_{t-1}$ and $x_{t}$. At last, the values of the vector and the regulated values are multiplied to be sent as an output and input to the next cell. The equation for the output gate is:
$o_{t}=\sigma(W_{o}[h_{t-1},x_{t}]+b_{o})$ (6)
For each time step t, the hidden state $h_{t}$ will be updated by the input of vector $x_{t}$, previous hidden cell state $h_{t-1}$ and previous cell state $c_{t-1}$. The vector formulas of pass forward along timesteps can be written as:
$h_{t}=o_{t}\odot tanh(c_{t})$ (7)
Where $W_{*}$ and $b_{*}$ denotes the weight matrix and bias of the corresponding gates respectively, $\sigma$() is the sigmoid activation function, $tanh$ is the hyperbolic tangent activation function, $\odot$ is the Hadamard product, which denotes an element-wise product, and $c_{t}$ is designed to accumulate the short memory.

### v2

@@ Abstract
There is still a lack of elaborate performance analysis at the lane-level due to costly data collection and difficulty in prediction modeling. Therefore, this study developed a multi-task deep learning approach to predict the lane-level pavement performance with a large amount of historical segment-level performance data. In specific, the prediction framework firstly employed an Long Short-Term Memory (LSTM) layer to capture the segment-level pavement deterioration pattern. Then multiple task-specific LSTM layers were designed based on number of lanes to capture lane-level differences in pavement performance. Finally, the multiple task-specific LSTM outputs were concatenated with auxiliary features for further training and the lane-level predictions were obtained after fully connected layer. The aforementioned prediction framework was validated with a real case in China. The proposed prediction framework also outperforms other ensemble learning and shallow machine learning methods in almost every lane.
@@
journal: Engineering Applications of Artificial Intelligence
@@
organization=School of Transportation,addressline=Southeast University, city=Nanjing, postcode=210000, state=Jiangsu, country=China
@@
Propose a deep learning approach for lane-level pavement performance prediction.
Introduce LSTM module to investigate temporal relations hidden in historical records.
Develop lane-level prediction module using multi-task learning approach.
Concatenate the deep learning module with a set of auxiliary roadway features.
Obtain good model performance under three regular roadway scenarios.
@@ 1 Introduction
The most widely used pavement performance measurements include pavement condition index (PCI), pavement quality index (PQI), and riding quality index (RQI). These measurements in a lane with shorter length can assist in identifying small but severely damaged road sections that are detrimental to traffic. The performance measurement prediction can further provide refined performance trends in future years for each lane. Utilizing the predicted lane-level performance values can further execute the preventive maintenance in a precisely and cost-effective way. Thus, lane-level performance predictions are significantly more useful for pavement maintenance than segment-level ones that merely provide one rough value for all lanes in a segment of 1 km. However, it is impossible to collect a huge amount of lane-level data in reality, despite the fact that many automated pavement detection methods have been proposed[1, 2], and it is also difficult to accurately predict lane-level performance in the next few years. To address such difficulty, this study is dedicated to developing a lane-level pavement performance prediction framework with a large amount of segment-level data.
@@
To address the difficulties during lane-level performance prediction, a multi-task deep learning approach is proposed for lane-level pavement performance prediction based on segment-level data to deal with those challenges. The model employs an LSTM-based shared layer to capture the common and representative changing pattern information from pavement performance time series. Based on the shared information, the task-specific heads are used to extract the details and changes related to a specific lane and generate predictions for the lane. Subsequently, the predictions are horizontally concatenated with auxiliary features, and the features distinguish roads with different decay patterns. At last, the task-specific output layer will handle the concatenation results to get the final lane-level predictions. To verify the proposed model, the real case was employed to check prediction accuracy of PCI, PQI, and RQI on three scenarios including one-way 2-lane, 3-lane, and 4-lane roads.
The remainder of this paper is organized as follows: In section 2, the related work on pavement performance prediction is summarized. In section 3, the structure of the proposed model is introduced detailed. Then, in section 4, the model is validated through comparative and ablation experiments based on a dataset that has been independently acquired. Finally, conclusions and guidelines for future work were presented.
@@ 2 Related work
The literature on pavement performance prediction models can be categorized into two typical groups of probabilistic reasoning models and shallow machine learning models. The group of probabilistic reasoning models primarily relies on specific mathematical equations or statistical laws. The group of shallow machine learning models contains typical data-driven models, rather than deriving from theoretical assumptions or prior knowledge.
2.1 Probabilistic reasoning models
@@
Probabilistic reasoning models predict pavement performance considering uncertainty and probability distribution of influencing factors. Specifically, the models can be further divided into empirical, mechanistic-empirical, and probabilistic models. Empirical models are based on traditional regression methods whose parameter are estimated with one dataset of pavement performance and other factors. Cao selected service life and cumulative number of axle loads as independent variables and applied cosine deterioration equations to predict PCI and rutting depth index of roads[5]. Wu examined the better performance of the S-shaped model on pavement performance prediction, comparing with the polynomial and exponential models[6]. Mechanistic-empirical models are usually based on mechanical principles and explore the relationship between pavement performance and pavement responses (i.e. stresses, strains, and deflections). Sidess combined a mechanistic-empirical approach with a regressive empirical approach to predict pavement performance [7]. Probabilistic models consider significant factors in pavement performance deterioration and describe the possible distribution or changing trend of pavement performance by statistical methods, such as Markov models and Monte Carlo simulations. Wang used the gray theory to predict pavement smoothness and used the fuzzy regression method to determine the coefficients for the gray prediction model[8]. Dong predicted asphalt pavement performance utilizing Markov models based on insufficient inspection data[9]. Wang put forward a Grey–Markov combination model to predict PCI[10]. Since the pavement deterioration rate may not be constant in the long term, Abaza developed a non-homogeneous Markov chain with different transition matrices[11]. Besides, Wasiq proposed a hybrid approach combining homogeneous and non-homogeneous Markov Chain to predict the PCI[12].
Probabilistic reasoning models perform well in particular with limited observation data, since such models assume underlying pavement deterioration pattern with classical distributions or equations, even without any proofs. The model performance mainly depends on very reasonable assumptions or statistically validated hypothesis. However, the pavement deterioration may be influenced by various factors including location-specific, lane-specific, time-specific, pavement design, and traffic. It is challenging in finding an universal pavement deterioration pattern applicable anywhere and anytime. Thus, the probabilistic reasoning models require a great deal of time and effort for deriving correct assumptions on pavement deterioration while predicting lane-level performance in a large-scale road network.
2.2 Machine learning models
Machine learning models are new trends in pavement performance prediction, typically including traditional machine learning algorithms and artificial neural networks (ANN). Such models are trained only with observation datasets, generally not introducing any assumptions or predefined equations/distributions. The trained modeling structure, as well as corresponding parameters, reveals the nonlinear relationships underlying the pavement deterioration. Regarding the traditional machine learning algorithms, Wang developed a hybrid grey relation analysis and support vector machine regression model to predict pavement performance[13]. Hu proposed a random forest-based prediction model and further enhanced model perforamnce with a grid search algorithm[8]. Ali investigated the combined effect of pavement distress on flexible pavement performance in two climate regions (wet freeze and wet no freeze) separated using multiple linear regression and artificial neural network[3]. Zhang and Kaloop both employed the Gaussian process regression model for predicting IRI. The distinction lies in that Zhang utilized a genetic algorithm to determine the optimal time series length[14], whereas Kaloop adopted regression modeling to select appropriate feature inputs[15]. In addition, the ensemble learning methods are also used to predict pavement performance due to their good performance in complicated industry problems, such as XGBoost[16] and ThunderGBM[17].
Regarding the deep learning based methods, the BPNN modeling framework was further improved with a differential evolution particle swarm optimization by Tao, while predicting PCI[18]. Cai proposed a causal-temporal graph convolution network, combining the graph convolution networks and LSTMs to simultaneously capture causal features and temporal features to predict pavement performance[19]. Bukharin proposed a two-stage model combining LSTM and ANN. The LSTM is used to learn the pavement deterioration pattern from time-series data, and the ANN further incorporates influencing factors for deterioration pattern[20]. Guo proposed a weighted multi-output neural network to predict IRI, faulting, longitudinal crack, and transverse crack in one unified framework[21]. Li proposed a model combining the BPNN and the LSTM to predict the highway pavement performance with a high precision of 100 m, based on ten-year traffic load data, climatic history, and maintenance records data[22]. Zhang developed a graph deep learning framework for network-based traffic prediction, and the frame models the dynamics of the traffic flow on a road network as a Markov chain on a directed graph. Still, it can not predict lane-level traffic volume which is important, especially in intersections.[23]. Yao used a deep reinforcement learning model to learn better maintenance strategies that maximize long-term cost-effectiveness in maintenance decision-making. In particular, each lane can have different treatments, and the long-term maintenance cost-effectiveness of the entire road is treated as the optimization goal. However, the lane-level data the model needs limits the wide use of the model[24].
The machine learning models are good at investigating and capturing hidden relationships in observation data, always yielding better performance once having enough big data. These models have been validated in the task of segment-level pavement performance prediction. However, the literature lacks examples of similar modeling structures for lane-level pavement performance prediction according to the authors’ investigation, in particular one unified prediction model resulting in pavement performance of multiple lanes. The task of lane-level pavement performance prediction is challenging due to much more complicated correlations and differences among lanes, compared to the task of segment-level prediction.
3 Method
@@ 3.1 Problem definition
This study is dedicated to proposing a lane-level pavement performance prediction model with both historical segment-level pavement performance data and auxiliary features influencing pavement deterioration.
$\mathbf{\hat{y}^{t+1}_{n,~{}j}}=f(\mathbf{y^{t-k:t}_{m}},~{}\mathbf{X_{m}}),~{}n=1,2,...,N.$ (1)
where,
$\mathbf{\hat{y}^{t+1}_{n,~{}j}}$ is the predicted pavement performance, like PCI, PQI, or RQI, on spatial prediction unit $j$ of lane $n$ at time step $t+1$;
$n$ is the lane number, ranged from 1 to $N$;
$N$ is the total number of lanes in one-way, usually 1, 2, or more;
$j$ is the spatial prediction unit number generally with a length of $100m$;
$t$ is the time step, generally representing the year;
$\mathbf{y^{t-k:t}_{m}}$ is the time series of pavement performance observed on spatial unit $m$ (i.e. one segment) from time step $t-k$ to time step $t$;
$\mathbf{X_{m}}$ is the auxiliary features measured on spatial unit $m$, in this study auxiliary features are time-invariant;
$m$ is the spatial measurement unit number, each unit is one road segment with a length of $1000m$, including multiple spatial prediction units $j$;
$f(\cdot)$ is the trained modeling structure.
@@ 3.2 Model overview
The section is focus on introducing the structure of the proposed multi-task deep learning model for lane-level pavement performance prediction, as shown in Figure 2. First, input features are processed with one segment-level deep learning module (a) that is based on LSTM, extracting general deterioration patterns at segment level from historical pavement performance data. Second, one task-specific deep learning module (b) yielding multiple heads for each lane-level prediction task and further learning underlying patterns with LSTM and fully connected neural network layers. In final, one concatenation module (c) horizontally mixes deep learning output with auxiliary features, resulting in final lane-level predictions by multiple fully connected neural networks. Regarding the model training, the mean square error (MSE) and their total sum are used to quantify the loss of each task and all tasks, respectively.
@@ 3.2.1 LSTM-based shared layer
The LSTM-based shared layer is initially used to handle the pavement performance time series at segment level. It is responsible for discovering and extracting common deterioration patterns of the whole road segment. LSTM is a recurrent neural network architecture designed to handle the vanishing and exploding gradient problems often encountered in traditional recurrent neural networks when dealing with long data sequences. The LSTM has a more complex internal structure than standard recurrent neural network cells. It consists of three main gates including the input gate, the forget gate, and the output gate. The input gate decides how much new information to let into the cell state. The forget gate determines which parts of the previous cell state to discard. The output gate controls how much of the cell state is output as the hidden state. The aforementioned gate mechanism allows LSTM to selectively remember or forget information, making it effective in long-term prediction tasks.
(1) Forget gate
The forget gate controls what information is removed from the memory cell. The gate takes two inputs of segment-level pavement performance $y^{tt}$ and hidden state at the previous time step $tt-1$. The both inputs are multiplied with one weight matrix $W_{g}$, which further adds one bias $b_{g}$, as in equation 2. The result is passed through an activation function that yields an output between 0 and 1. In specific, if the output is 0, the information is forgotten. In contrast, if the output is 1, the information is retained for further computation in future time steps.
$g_{tt}=\sigma(W_{g}[h_{tt-1},y^{tt}]+b_{g})$ (2)
where,
$y^{tt}$ is set of pavement performance on road segment $1$ to $M$ at time step $tt$, $\left[y^{tt}_{1},y^{tt}_{2},\ldots,y^{tt}_{M}\right]$;
$g_{tt}$ is the forget gate information indicator at time step $tt$;
$\sigma(\cdot)$ is the sigmoid activation function;
$W_{g}$ represents the weight matrix used in the forget gate;
$h_{tt-1}$ is the hidden state at the previous time step $tt-1$, integrating information from the first time step to the time step $tt-1$;
$b_{g}$ is the bias in the forget gate.
(2) Input gate
The input gate controls what information is added to each memory cell. On one hand, the information indicator $i_{tt}$ is initially computed with weighted inputs and sigmoid function, similar as computation in forget gate. On the other hand, the initial cell state $\stackrel{{\scriptstyle\sim}}{{c_{tt}}}$ is also computed with weighted inputs, but is activated with the tanh function, shown in equation 4. Such activation function will result in an output ranging from -1 to 1. The cell state $c_{tt}$ is finally obtained with equation 5, combining together cell state at previous time step, the forget gate output at current time step, the new information in input gate, and the initial cell state.
$i_{tt}=\sigma(W_{i}[h_{tt-1},y^{tt}]+b_{i})$ (3)
@@
$c_{tt}=g_{tt}\odot c_{tt-1}+i_{tt}\odot\stackrel{{\scriptstyle\sim}}{{c_{tt}}}$ (5)
where,
$i_{tt}$ is information indicator in the input gate at time step $tt$;
$W_{i}$ is the weight matrix in the input gate;
$b_{i}$ is the bias in the input gate;
$\stackrel{{\scriptstyle\sim}}{{c_{tt}}}$ is the initial cell state at time step $tt$;
$\tanh(\cdot)$ is the tanh activation function;
$W_{c}$ is the weight matrix in the memory cell;
$b_{c}$ is the bias in the memory cell;
$c_{tt}$ is the cell state at time step $tt$;
$c_{tt-1}$ is the cell state at time step $tt-1$;
$\odot$ is the Hadamard product denoting an element-wise product.
@@
The output gate controls what information is output from the memory cell. First, the output information indicator $o_{tt}$ is generated through activating weighted inputs as other information indicators. Then, the hidden state $h_{tt}$ will be updated by the output information indicator and the tanh activation of cell state, as follows.
$o_{tt}=\sigma(W_{o}[h_{tt-1},y^{tt}]+b_{o})$ (6)
$h_{tt}=o_{tt}\odot tanh(c_{tt})$ (7)
where,
$o_{tt}$ is the output information indicator in the output gate at time step $tt$;
$W_{o}$ represents the weight matrix in the output gate;
$b_{o}$ is the bias in the output gate;
$h_{tt}$ is the hidden state at time step $tt$.